A solution of hydrogen peroxide is 13.7% by mass. What is the molarity of the
solution? Assume that the solution has a density of 1.01 g/mL.
1
Expert's answer
2021-09-17T02:08:29-0400

13.7g "H_2O_2" in 100g solution

D=1.01 g/ml

"volume=\frac{100}{1.01}=99 \space ml"

moles of solute = "\frac{13.7}{34}=0.4"

"molarity=\frac{moles \space of\space solute} {v(L)}\\=\frac{0.4}{0.099}\\=4.04"